OpenStudy (anonymous):

Solve. 14x2 + 42x + 70 = 0

OpenStudy (anonymous):

I had gotten x = -3 +/- i√11 over 2 and I just wanted to check my answer. :)

OpenStudy (ajprincess):

Divide the whole equation by 14. then u will get x^2+3x+5=0 x^2+3x+9/4=-5+9/4 (x+3/2)^2=(-20+9)/4 (x+3/2)^2=-11/4 x+3/2=+/-sqrt(-11)/2 x=(-3+/-11i)/2

OpenStudy (ajprincess):

ur answer is right.

OpenStudy (anonymous):

Thank you!

OpenStudy (ajprincess):

yw